OpenStudy (anonymous):

I have this question on my homework and I need help solving it. I do have part of it done, but I need to improvise as far as expressing the problem goes. Consider the area under one arch of the curve y(t) = ksin(wt) for t ≥ 0 where k and w are positive constants. Note: Use lower case k and w when needed. (a) Using the form (Integral from A to B of ksin(wt) dt, find A and B. Part b asked to find u and du, which I got on my own: u = wt and du = w dt (c) What are the new endpoints for the integral in terms of u? I'm pretty sure I can figure out part d, which is asking for the exact area.

jimthompson5910 (jim_thompson5910):

you need to solve y(t) = 0 y(t) = ksin(wt) 0 = ksin(wt) ksin(wt) = 0 k = 0 or sin(wt) = 0 let's assume k is nonzero, that would have to mean that sin(wt) = 0 what is t if sin(wt) = 0 ?

OpenStudy (anonymous):

I've done a bit more work since I posted this earlier. I'm on the last couple parts now, c and d. A I found to be zero, since it says t ≥ 0. Then I found B to be pi/w. So, substituting in u and du, A stays 0, but what is B?

jimthompson5910 (jim_thompson5910):

u = w*t if t = 0 (first root where t >= 0), then u = w*t u = w*0 u = 0 so A = 0 if t = pi/w (second root after t = 0) u = w*t u = w*pi/w u = pi and B = pi

OpenStudy (anonymous):

Alright, so let's see... k is pulled out of that equation since it is a constant, I'm assuming. So that turns out looking like this? \[k \int\limits_{0}^{\pi}\sin(u)du\] Sorry, I'm not that great at calculus, so I'm probably going to be coming here for help a lot in the future.

jimthompson5910 (jim_thompson5910):

you're missing one thing though

jimthompson5910 (jim_thompson5910):

u = w*t du = w*dt ----> dt = du/w so you should have du/w and not just du

jimthompson5910 (jim_thompson5910):

what you can do is think of du/w as du*(1/w) then pull out the constant 1/w

jimthompson5910 (jim_thompson5910):

you should really have this integral \[\Large \frac{k}{w} \int_{0}^{\pi} \sin(u)du\]

OpenStudy (anonymous):

Oh, I was closer than I thought I'd be. According to the new equation, that results in something like this? \[k/w * (1-\cos(\pi*t))/t\] I got the (1-cos(pi*t))/t as the result of the integral without the k/w.

jimthompson5910 (jim_thompson5910):

don't forget about the limits of integration

OpenStudy (anonymous):

The limits? You mean from 0 to pi?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

Let \[\LARGE g(u) = \int \sin(u)du\] you should find that \[\LARGE g(u) = -cos(u) + C\] to compute \[\LARGE \frac{k}{w}\int_{0}^{\pi} \sin(u)du\], you'll need to compute \[\LARGE \frac{k}{w}*\left(g(\pi) - g(0)\right)\]

OpenStudy (anonymous):

g(pi) I get to be 1, and g(0) I get to be -1, so g(pi) - g(0) is 2, correct? So that comes out to 2k/w.

jimthompson5910 (jim_thompson5910):

yep, \[\LARGE \frac{k}{w}\int_{0}^{\pi} \sin(u)du=\frac{2k}{w}\] is correct

OpenStudy (anonymous):

Weird, I kept getting something else for that one part of the equation for some odd reason. But, I ended up with the right answer, so I'll go back and look through this once I've got some free time on my hands after classes tomorrow. I still have a couple more problems for this homework assignment that I need help on. They have multiple parts, but it is the last two problems of my homework. Should I start a new thread?

jimthompson5910 (jim_thompson5910):

yes it's always a good idea to start a new post for a new problem (to avoid clutter and lag)

OpenStudy (anonymous):

Alright, thank you very much for the help. And your patience. :)

jimthompson5910 (jim_thompson5910):

you're welcome